OpenStudy (anonymous):

a simple dart board has 3 areas...the main board has a radius of 9 in, there is a circle with a radius of 6 in, and the bullseye has a radius of 2 in. what is the probability of a random dart landing inside the bullseye. I'm a little confused.

OpenStudy (anonymous):

area of the bulls eye divided by total area

OpenStudy (anonymous):

\[\frac{4\pi}{81\pi}=\frac{4}{81}\]

OpenStudy (anonymous):

as the area of a circle with radius \(r\) is \(\pi r^2\)

OpenStudy (anonymous):

ok so the probability would be 4/81 as well

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw